SFG SPECTROSCOPY OF FUNCTIONALIZED POLY (p-XYLYLENE) INTERFACES

\underline{ARTHUR McCLELLAND}, HSIEN-YEH CHEN, JOERG LAHANN, ZHAN CHEN, UNIVERSITY OF MICHIGAN, ANN ARBOR, MI.

Poly (p-xylylene), also known as parylene, is an important polymer in many applications from sealants on electronics to biomedical implants. By adding different functional groups, the surface properties can be tailored for specific applications. SFG spectroscopy allows us to probe the surface functional groups in situ in aqueous environments and at buried interfaces.